Calculate the volume, in litres, occupied by 1.0 mol of an ideal gas at STP.


1
Expert's answer
2021-10-26T05:34:10-0400

Ideal gas law

pV=nRT

R = 0.08206 L×atm/mol×K

At STP

p = 1 atm

T= 273 K

"V = \frac{nRT}{p} \\

V= \frac{1.0 \times 0.08206 \times 273}{1} = 22.4 \;L"

Answer: 22.4 L